OpenStudy (anonymous):

x(2)^+x-6=0 work shown

OpenStudy (mathmate):

x^2+x-6=0 x^2-2x +3x-6=0 x(x-2) + 3(x-2)=0 Can you extract the common factor and continue?

OpenStudy (anonymous):

i need work shown step by step please

OpenStudy (mathmate):

I have supplied the first two steps to factorize the given polynomial. You need to complete the work by extracting the common factor (x-2) and write down the solutions.